# Splittable Spanning Trees and Balanced Forests in Dense Random Graphs

### v1

@@ Abstract.
Weighted equitable partitioning of a graph $G$ has been of interest lately due to several applications, including redistricting, network algorithms, and image decomposition. Equitably partitioning a graph according to the spanning-tree metric, which assigns a weight to each component equal to its number of spanning trees, and overall weight equal to the product of the weights of its components, has been of mathematical and practical interest because it typically favors partitions with more compact pieces. An appealing algorithm suggested by Charikar et al. is to sample a random spanning tree and remove $k-1$ edges, producing a random forest. If the components of the forest form a balanced partition, the partition is equitable under an easily computed acceptance probability. Cannon et al. recently showed that spanning trees on grid graphs and grid-like graphs on $n$ vertices are splittable into $k$ equal sized pieces with probability at least $n^{-2k}$, leading to the first rigorous sampling algorithm for a class of graphs. We present complementary results showing that spanning trees on dense random graphs also have inverse polynomial probability of being splittable, specifically with probability at least $n^{(-k/2)}$ for both the $G(n,p)$ and $G(n,m)$ models, giving another class of graphs where equitable partitions can be efficiently sampled exactly. Interestingly, $G(n,m)$ is the easier case, and we infer the property for random graphs from $G(n,p)$ indirectly. These proofs also guarantee that the up-down walk for almost uniformly sampling forests without restrictions on the sizes will produce a balanced forest with inverse polynomial probability, giving another provably efficient randomized method for generating equitable partitions.
Further, we show that problems with the well-studied ReCom algorithm for equitable partitioning are more extensive than previously known, even in special cases that were believed to be more promising. We present a family of graphs where the Markov chain fails to be irreducible when it must keep the components perfectly equitable; yet when the chain is allowed an imbalance of just one vertex between components, the rejection sampling step may take exponential time. Thus, ReCom, which combines just two components at each step, can suffer from splittability as much as the direct algorithm that chooses a spanning tree and removes $k-1$ edges at once. This is true even when the graph satisfies desirable properties that have been conjectured to be sufficient for fast sampling, such as Hamiltonicity, having low degree, or being an induced subgraph of the triangular lattice.
@@ 1. Introduction
Weighted, equitable partitions of a graph have garnered a lot of interest lately, especially in the context of political redistricting. Given an integer $k>1$ and a graph $G=(V,E)$ with $|V|=kn$, we are interested in sampling all partitions of the vertices of $G$ into $k$ connected components $V=(V_{1},\dots,V_{k})$ with probability proportional to $\Pi_{i=1}^{k}w(V_{i})$ for some weight function $w:V^{\prime}\rightarrow\mathbb{R}$ for any $V^{\prime}\subseteq V$. A popular distribution is based on the spanning tree weight, which sets $w(V^{\prime})$ to be the number of spanning trees of $V^{\prime}$. Notice that sampling partitions according to the spanning tree weight is equivalent to sampling balanced forests, where each connected component of the forest has exactly $n$ vertices.
@@
In the context of redistricting, random weighted equitable partitions are sampled to form an ensemble against which proposed maps can be tested for political gerrymandering. Each vertex of the graph represents a small indivisible region, or precinct, and the components of a partition represent the districts from which representatives are elected. This methodology has been used for court cases and various amicus briefs [11, 19, 6]. While the requirements for acceptable political maps vary state to state, mathematicians have generally agreed that the spanning tree weight is a desirable baseline because spanning trees are easy to generate and they typically produce maps with compact districts [16]; when other weightings are preferable, parallel tempering has been successfully used to interpolate between the spanning tree weight and other metrics [6, 5]. Uniformly generating balanced forests is also desirable in the contexts of clustering in biological and social networks [7, 15] and image decomposition [13].
Algorithms for sampling balanced forests:  The most widely used algorithm for sampling balanced forests is ReCom (for recombination) [12] and its variants, which provably converge to the spanning tree weight [9, 6] [^1] For simplicity of exposition, we refer to this group of algorithms simply as “ReCom,” although our results rely on any of the modified versions which have provably correct stationary distributions [9, 6]. These have been particularly popular among mathematicians assisting politicians and lawyers by establishing a baseline against which they can identify outliers. The algorithms all use Markov chains that walk on balanced forests (or weighted partitions) by iteratively joining two adjacent pieces and then splitting this joint region into two equal sized pieces with the correct conditional probabilities. When a tree cannot be equitably split by removing a single edge, the move is rejected and the Markov chain proceeds to propose a new recombination move.
For this approach to be efficient and reliable, it is necessary that the recombination move connects the state space of conditions, the chain is rapidly mixing, and subgraphs of $G$ are splittable with at least an inverse polynomial probability. Unfortunately, it has been shown that there are potentially issues with each of these three requirements. The chain is not guaranteed to mix quickly, and indeed there exist graphs where it mixes slowly (double cycle graphs) [11]. Worse, the state space may not even be connected; a cycle with $3n$ vertices can be partitioned into $3$ pieces of size $n$ along the cycle in $n$ different ways, but no moves are possible via ReCom aside from relabeling the pieces. In fact, no graphs except the complete graph are known to be connected under perfectly balanced ReCom moves [8]. More strikingly, the problem of determining connectivity of the state space under ReCom moves is PSPACE-complete, even for planar graphs [3]. This means that when running this algorithm on a given graph, the user has no way of knowing if the results are close to stationarity, and so we cannot trust this result for the purpose of sampling without additional proofs about some restricted classes of graphs.
Natural modifications to ReCom have been proposed to increase its connectivity and potentially its mixing rate by allowing some slack on the piece sizes. However, even with this allowance, it still has only been shown to be connected in special cases, such as the double cycle graph [11], a triangular region of the triangular lattice with slack at most 1 [8], and Hamiltonian graphs with a significantly larger linear-sized slack [3]. For most graphs little is known about the probability of a random spanning tree being “splittable” into two (or more) equal sized pieces.
Alternative approaches for directly generating balanced forests also have been proposed, relying on a random spanning tree’s splittability into $k$ pieces, defined below.
@@
A tree is $k$-splittable if there exist $k-1$ edges that, when removed, create a graph whose $k$ connected components all have the same number of vertices.
@@
The simplest alternative sampling algorithm generates uniform spanning trees on all of $G$ (using, for example, the loop erased random walk algorithm of Wilson [18]) and checks whether each tree is $k$-splittable, which can be done in efficiently if $k$ is constant. Note that any splittable tree has a unique set of $k-1$ edges whose removal leaves a balanced forest, so rejection sampling will generate such forests uniformly. If the tree is not splittable, it is resampled. This sampling algorithm produces samples from exactly the right distribution, but the running time depends on the rejection rate. A second algorithm uses a Markov chain to generate samples from approximately the right distribution, also using rejection sampling.
@@
Approximate Markov chain algorithms also exist in this setting, using the up-down walk [11] to efficiently sample random forests with exactly $k$ components but with no restriction on the sizes of the components. This chain has a $O(n^{2}\log(n))$ mixing time on all graphs and is known to be ergodic. Balanced forests will be generated almost uniformly and all others will be rejected and resampled, so the algorithm will be fast if the rejection rate is not too large. Note that the probability that a random tree is splittable is polynomially related to the likelihood a random forest is balanced (see Lemma 2.2), so either both or neither of these algorithms will be efficient.
@@
Charikar et al. [11] conjectured that many graphs are splittable with inverse polynomial probability, including rectangular regions of $\mathbb{Z}^{2}$. Recently, Cannon, Pegden and Tucker-Foltz [10] gave a stunning proof of this conjecture for planar grid graphs and grid-like graphs. However, planarity is critical to their argument because they rely on the dual of the loop erased random walk of Wilson to bound the likelihood of specific families of forests being generated to show that these already occur with at least $1/n^{2}$ probability, suppressing dependencies on $k$. Little is known about the splittability of any other classes of graphs.
Results:  When $k=2$, ReCom is equivalent to generating and splitting a tree of the entire graph: ReCom only makes a move when it finds a $2$-splittable random spanning tree of the entire graph $G$, and when it does, it immediately moves to a balanced $2$-partition selected according to the spanning tree weight. Allowing some small slack on the piece sizes is a potentially necessary modification to guarantee connectivity of the state space [3, 8], or to avoid slow mixing [11]. We present a family of graphs where ReCom with strictly balanced piece sizes cannot make any move, and yet allowing even a single vertex of slack on the sizes of two pieces makes it exponentially unlikely that ReCom will generate a balanced partition. These graphs have effectively every relevant property that has been conjectured are sufficient to avoid such behavior, including Hamiltonicity [3], are induced subgraphs of the triangular lattice [8], implying that they are also planar graphs with maximum degree bounded by a constant.
Theorem 1.
There exists a family of graphs where the total spanning tree weight of partitions into $3$ pieces of sizes $\{n-1,n,n+1\}$ is exponentially higher than the total weight of partitions into $3$ pieces of size $n$. Further, there exist such graphs that are induced Hamiltonian subgraphs of the triangular lattice, and have a disconnected state space when using ReCom without slack.
Next, to examine the mixing time of ReCom on the complete graph $K_{kn}$, we can consider a “bin-mixing” Markov chain on $k$ bins of $n$ balls each, where a move consists of combining $2$ bins and randomly mixing the balls between them such that each bin ends up with $n$ balls. We see that when running ReCom on the complete graph, each move where it finds a $2$-splittable spanning tree is equivalent to a move on the bin-mixing Markov chain, and each one where it does not is a move where it does nothing. ReCom on this graph can be considered as a lazy version of the bin-mixing Markov chain. If the bin-mixing Markov chain mixes in time $M$, then ReCom mixes in time $M~{}\times$ the expected number of samples of spanning trees of $K_{kn}$ needed to find a $k$-splittable one. We therefore see that the mixing time of ReCom on this graph is proportional to the inverse probability that a random tree is splittable. To estimate this probability, we prove the following theorem.
@@
Finally, we prove that a polynomial fraction of the set of k-component forests on dense regular graphs will be balanced, for any constant k. This complements the work of Cannon et al. [10], who established a similar result for the m x n grid and a related class of “grid-like” graphs.
@@
Note that Cannon et al. [10] showed the proportion of spanning trees that are splittable on the $\sqrt{n}\times\sqrt{n}$ grid scales by at least $n^{2-2k}$, so these bounds on splittability for dense graphs are much more favorable.
@@
What makes analysis of the splittability of dense graphs especially challenging, and even somewhat surprising, is that spanning trees that appear with higher probability than others are less likely to be splittable. While any two spanning trees are equally likely to appear in a random graph, one tree may be more likely to appear in graphs that contain fewer spanning trees, thus making it more likely to be selected at random. For example, a star with three leaves is more likely than a path of three edges to be selected as a random spanning tree of a random graph on four vertices. Alon et al. [4] have shown that the diameter of random spanning trees of dense random graphs grow as $\sqrt{n}$, as was previously shown by Szekeres [17] for the complete graph. Trees with smaller diameter would seem less likely to be splittable than those with large diameter, so inverse polynomial bounds on the splittability of uniform spanning trees on dense graphs is more suprising than on sparse graphs such as the grid, where the diameter of a uniform spanning tree is linear.
Another difficulty is that balanced forests tend not to be the mode of a histogram showing the probabilities of splits of various sizes, even when $k=2$. In fact, when on the complete graph, balanced splits contain the fewest forests out of all possible partition sizes, as shown in Figure 2. This means that standard techniques showing that this distribution is log-concave or unimodal will certainly fail and the calculations need to be more precise.
@@ 2.1. The spanning tree weight and splittability
Given any tree $T(V,E)$ with $kn$ vertices, for some integer $k$, there exists at most one set $F\subseteq E$ of $k-1$ edges such that the removal of $F$ from $T$ creates $k$ connected components, each with size $n$.
@@
When $k=1$, clearly $F$ can only be the empty set. Assume this lemma holds for splits into at most $k-1$ pieces and proceed by induction.
If there exist two sets $F_{1},F_{2}$ satisfying the conditions of the Lemma, consider the graph created by contracting the components of $T$ that would result from removing $F_{1}$. This is a tree with $k$ vertices, and therefore has a leaf, corresponding to a subtree $S$ of $T$ with $n$ vertices and exactly one edge $e$ between it and $V-S$. $F_{2}$ cannot contain any edge in $S$, as if it did, it would split $S$ into multiple connected components in $T-F_{2}$, only one of which can be connected to $e$, and the other of which would be a proper subset of $S$. A proper subset of $S$ has fewer than $n$ vertices, which is impossible, so $F_{2}$ cannot contain any edge in $S$.
If $F_{2}$ contains no edges in $S$, all of $S$ is connected in $T-F_{2}$. Then to avoid the connected component containing $S$ having more than $n$ vertices, $F_{2}$ must contain $e$, as otherwise, $S$ is connected and connected to $e$ in $T-F_{2}$. Then both $F_{1}$ and $F_{2}$ contain $e$, and removing $e$ splits $T$ into a connected component with $n$ vertices and one with $(k-1)n$ vertices. Then both $F_{1}$ and $F_{2}$ contain subsets of $k-2$ edges splitting this other subgraph into pieces of size $n$, so by induction, they contain the same edges. Then $F$ is unique. ∎
@@
We note that the probability that a random spanning tree is balanced is the ratio of the number of $k$-splittable spanning trees to the number of spanning trees of a graph, and that the probability that a random k-partition weighted by the spanning tree weight is balanced is the ratio of the total weight of all balanced $k$-partitions to the total weight of all $k$-partitions. Then this is a direct consequence of the next two lemmas. A similar result is proven in [10]. ∎
@@
We see that each such forest can, by adding $k-1$ edges of the components, be made into a spanning tree, and since every spanning tree can be split into $k$ pieces by arbitrarily removing $k-1$ edges, every spanning tree can be made this way. Since there are at most $n^{2}$ edges in $G$, the number of spanning trees is at most $n^{2(k-1)}$ times the number of such forests. Similarly, every such forest can be made by removing $k-1$ edges from a spanning tree, the number of such forests is at most $n^{2(k-1)}$ times the number of spanning trees (this bound can be reduced to $n^{k-1}$ by the structure of the tree). ∎
@@
By Lemma 2.1, each splittable tree uniquely defines a $k$-balanced spanning forest, so the number of them is at most the number of spanning trees. Further, each balanced tree can be defined by a balanced forest and $k-1$ edges, so there are at most $n^{2(k-1)}$ times as many spanning trees as balanced forests. ∎

### v2

@@ Abstract.
We consider the probability that a spanning tree chosen uniformly at random from a graph can be partitioned into a fixed number $k$ of trees of equal size with the removal of k1k-1 edges. In that case, the spanning tree is called splittable. Splittable spanning trees are useful in algorithms for sampling balanced forests, forests whose components are of equal size, and for sampling partitions of a graph into components of equal size, with applications in redistricting, network algorithms, and image decomposition. Cannon et al. recently showed that spanning trees on grid graphs and grid-like graphs on $n$ vertices are splittable into $k$ equal sized components with probability at least $n^{-2k}$, leading to the first rigorous sampling algorithm for balanced forests for any class of graphs. Focusing on the complementary case of dense random graphs, we show that random spanning trees again have inverse polynomial probability of being splittable; specifically, a random spanning tree is splittable with probability at least $n^{(-k/2)}$ for both the $G(n,p)$ and $G(n,m)$ models when $p=\Omega(1/\log n)$, giving the first dense class of graphs where partitions of equal size can be sampled efficiently.
In addition, we present an infinite family of graphs with properties that have been conjectured to ensure splittability (i.e., Hamiltonian subgraphs of the triangular lattice) and prove that random spanning trees are not splittable with more than an exponentially small probability. As a consequence of this, we show that a family of widely-used Markov chain algorithms for sampling partitions of equal size will fail on this family of graphs if their state spaces are restricted to equal-size partitions. Moreover, we show these algorithms will be inefficient if their state spaces are generalized to include any unbalanced partitions, suggesting barriers for sampling balanced partitions in other classes of sparse graphs.
@@ 1. Introduction
Sampling balanced forests of a graph has garnered a lot of interest lately, especially in the context of political redistricting. Given an integer $k>1$ and a graph $G=(V,E)$ with $|V|=kn$, the problem is to sample uniformly from the set of forests of $G$ consisting of $k$ trees, each of size $n$. Sampling balanced forests is equivalent to sampling partitions of the vertices of $G$ into $k$ connected components $V=(V_{1},\dots,V_{k})$ with probability proportional to $\Pi_{i=1}^{k}w(V_{i})$, where the weight $w(V_{i})$ is the number of spanning trees of $V_{i}$. This is known as the spanning tree weight.
@@
When used for redistricting, random weighted equitable partitions are sampled to form an ensemble against which proposed maps can be tested for political gerrymandering. Each vertex of the graph represents a small indivisible region, or precinct, and the components of a partition represent the districts from which representatives are elected. This methodology has been used for court cases and various amicus briefs [charikar2022complexity, zhao2022, DukeReCom1]. While the requirements for acceptable political maps vary state to state, mathematicians have generally agreed that the spanning tree weight is a desirable baseline because spanning trees are easy to generate and they typically produce maps with compact districts [procacciaTuckerFoltz2021compactness]; moreover, when other weightings are preferable, parallel tempering has been successfully used to interpolate between the spanning tree weight and other useful metrics [DukeReCom1, DukeReCom2]. Uniformly generating balanced forests is also desirable in the contexts of clustering in biological and social networks [camilus2012, junker] and image decomposition [fortunato2010].
1.1. Algorithms for sampling balanced forests
There are three families of algorithms for sampling balanced forests, all of which rely on a random spanning tree’s “splittability” into $k$ pieces, which we now define.
@@ 1.1. Algorithms for sampling balanced forests
A tree is $k$-splittable if there exist k1k-1 edges that, when removed, create a graph whose $k$ connected components all have the same number of vertices.
The most widely used algorithm for sampling balanced forests is ReCom (for recombination) [recom] and its variants, which provably converge to the distribution defined by spanning tree weight [revrecom, DukeReCom1] [^1] For simplicity of exposition, we refer to this group of algorithms simply as “ReCom,” although our results rely on any of the modified versions which have provably correct stationary distributions [revrecom, DukeReCom1]. ReCom uses Markov chains that walk on balanced forests (or weighted partitions) by iteratively choosing two adjacent pieces, finding a spanning tree on the vertices in the union of these piece, and then partitioning this joint region into two equal sized pieces by splitting the spanning tree in two, if possible. When a tree cannot be equitably split by removing a single edge, the move is rejected and the Markov chain proceeds to propose a new recombination move.
For this approach to be efficient and reliable, it is necessary that the recombination move connects the state space of partitions, the chain is rapidly mixing, and subgraphs of $G$ contain splittable spanning trees with at least an inverse polynomial probability. Unfortunately, it has been shown that there are potentially issues with each of these three requirements. In particular, no graphs except the complete graph are known to be connected under perfectly balanced ReCom moves. Moreover, Charikar et al. gave a family of graphs where strictly following ReCom moves will be slowly mixing [charikar2022complexity]. Natural modifications to ReCom have been proposed to increase its connectivity and potentially its mixing rate by allowing some slack on the piece sizes. Yet, even with this allowance, the state space has only been shown to be connected in special cases [charikar2022complexity, cannontri, akitaya2022reconfiguration_recombination].
@@
Another family of sampling algorithms work by generating uniform spanning trees on all of $G$ (using, for example, the loop erased random walk (LERW) algorithm of Wilson [wilson]) and checking whether each tree is $k$-splittable, which can be done efficiently if $k$ is constant. Note that any splittable tree has a unique set of k1k-1 edges whose removal leaves a balanced forest (see Lemma 2.1), so rejection sampling will generate such forests uniformly. If the tree is not splittable, it is resampled. This algorithm produces samples from exactly the right conditional distribution, but the running time depends on the rejection rate.
@@
Markov chains that walk on the set of (not necessarily balanced) forests with $k$ components also exist, such as the up-down walk [charikar2022complexity]. This chain has a $O(n^{2}\log(n))$ mixing time on all graphs and is known to be ergodic, but again we need to employ rejection sampling whenever the forests are imbalanced. The remaining balanced forests will be generated almost uniformly and all other forests will be rejected and resampled, so the overall algorithm will be fast only if the rejection rate is not too large. Note that the probability that a random tree is splittable is polynomially related to the likelihood a random forest is balanced (see Lemma 2.2), so either both the LERW and up/down walk approaches to sampling balanced forests are efficient or neither will be.
@@
As all of these algorithms rely on the probability that random spanning trees are splittable, it is worthwhile to characterize which classes of graphs have this property. Charikar et al. [charikar2022complexity] conjectured that many graphs are splittable with inverse polynomial probability, including rectangular regions of $\mathbb{Z}^{2}$. Recently, Cannon, Pegden and Tucker-Foltz [cannon2024grid] gave a stunning proof of this conjecture for planar grid graphs and grid-like graphs. Planarity is critical to their argument because they rely on the dual of the loop erased random walk to bound the likelihood of specific families of forests being generated to show that these already occur with at least $1/n^{2}$ probability, suppressing dependencies on $k$.
Little is known about sufficient conditions to ensure splittability in other classes of graphs. Akitaya et. al. [akitaya2022reconfiguration_recombination] have conjectured that Hamiltonicity is sufficient, and Cannon et. al. [cannontri] have conjectured that it is sufficient for a graph to be an induced subgraph of the triangular lattice. Little is known as well about sufficient conditions to ensure fast mixing of ReCom. It has been conjectured that it is necessary to allow slack on the piece sizes to guarantee connectivity of the state space [akitaya2022reconfiguration_recombination, cannontri], or to avoid slow mixing [charikar2022complexity].
1.2. Results
While it is tempting to believe the conventional wisdom that there are simple conditions on graphs that are sufficient to ensure their splittability with high probability, such as Hamiltonicity, sparsity, and planarity (e.g., [akitaya2022reconfiguration_recombination, cannontri, charikar2022complexity]), we give counterexamples to each of these conjectures. In particular, in Theorem 3, Section 3, we present an infinite family of graphs on $3n$ vertices that are induced Hamiltonian subgraphs on the triangular lattice, where the only equitable partitions of the graph into three connected components are the $n$ cyclical partitions of the Hamiltonian cycle at i,i+n,i+2n,i,\ i+n,\ i+2n, for $0\leq i\leq n-1$. However, the probability of a random spanning tree being splittable on this family is exponentially small. This implies that all of the algorithms for generating balanced partitions mentioned above, including ReCom, LERWs, and the up-down walk, all provably require exponential time.
This infinite family of graphs on $3n$ vertices also constitutes a troubling counterexample for ReCom with slack. In particular, if the state space for ReCom requires the subgraphs to be perfectly balanced, then the state space is disconnected; however, if even the smallest amount of slack is permitted and the state space includes partitions into pieces of size $\{n-1,n,n+1\}$, then ReCom will require exponential time to find a balanced partition. This example gives strong evidence that the hope to characterize which sparse graphs are splittable will be more challenging than previously believed.
[…]
We have the following theorem.
@@ 1.2. Results
Finally, we prove that a polynomial fraction of the set of k-component forests on dense regular graphs will be balanced, for any constant k. This complements the work of Cannon et al. [cannon2024grid], who established a similar result for the m x n grid and a related class of “grid-like” graphs.
@@
Note that Cannon et al. [cannon2024grid] showed the proportion of spanning trees that are splittable on the $\sqrt{n}\times\sqrt{n}$ grid scales by at least $n^{2-2k}$, so these bounds on splittability for dense graphs are much more favorable.
@@
What makes analysis of the splittability of dense graphs especially challenging, and even somewhat surprising, is that spanning trees that appear with higher probability than others are less likely to be splittable. While any two spanning trees are equally likely to appear in a random graph, one tree may be more likely to appear in graphs that contain fewer spanning trees, thus making it more likely to be selected at random. For example, a star with three leaves is more likely than a path of three edges to be selected as a random spanning tree of a random graph on four vertices. Alon et al. [alon20] have shown that the diameter of random spanning trees of dense random graphs grow as $\sqrt{n}$, as was previously shown by Szekeres [szekeres83] for the complete graph. Trees with smaller diameter would seem less likely to be splittable than those with large diameter, so inverse polynomial bounds on the splittability of uniform spanning trees on dense graphs is more suprising than on sparse graphs such as the grid, where the diameter of a uniform spanning tree is linear.
Another difficulty is that balanced forests tend not to be the mode of a histogram showing the probabilities of splits of various sizes, even when $k=2$. In fact, when on the complete graph, balanced splits contain the fewest forests out of all possible partition sizes. (See Figure 2 in Section 4.) This means that standard techniques showing that this distribution is log-concave or unimodal will certainly fail and the calculations need to be more precise.
@@ 2.1. The spanning tree weight and splittability
The spanning tree weight is often used to generate equitable partitions of a graph because the conditional distribution on balanced forests is uniform.
@@
Given any tree $T(V,E)$ with $kn$ vertices, for some integer $k$, there exists at most one set $F\subseteq E$ of k1k-1 edges such that the removal of $F$ from $T$ creates $k$ connected components, each with size $n$.
@@
When $k=1$, clearly $F$ can only be the empty set. Assume this lemma holds for splits into at most k1k-1 pieces and proceed by induction.
If there exist two sets $F_{1},F_{2}$ satisfying the conditions of the Lemma, consider the graph created by contracting the components of $T$ that would result from removing $F_{1}$. This is a tree with $k$ vertices, and therefore has a leaf, corresponding to a subtree $S$ of $T$ with $n$ vertices and exactly one edge $e$ between it and VSV-S. $F_{2}$ cannot contain any edge in $S$, as if it did, it would split $S$ into multiple connected components in $T-F_{2}$, only one of which can be connected to $e$, and the other of which would be a proper subset of $S$. A proper subset of $S$ has fewer than $n$ vertices, which is impossible, so $F_{2}$ cannot contain any edge in $S$.
If $F_{2}$ contains no edges in $S$, all of $S$ is connected in $T-F_{2}$. Then to avoid the connected component containing $S$ having more than $n$ vertices, $F_{2}$ must contain $e$, as otherwise, $S$ is connected and connected to $e$ in $T-F_{2}$. Then both $F_{1}$ and $F_{2}$ contain $e$, and removing $e$ splits $T$ into a connected component with $n$ vertices and one with (k1)n(k-1)n vertices. Then both $F_{1}$ and $F_{2}$ contain subsets of k2k-2 edges splitting this other subgraph into pieces of size $n$, so by induction, they contain the same edges. Then $F$ is unique. ∎
@@
We note that the probability that a random spanning tree is balanced is the ratio of the number of $k$-splittable spanning trees to the number of spanning trees of a graph, and that the probability that a random k-partition weighted by the spanning tree weight is balanced is the ratio of the total weight of all balanced $k$-partitions to the total weight of all $k$-partitions. Then this is a direct consequence of the next two lemmas. A similar result is proven in [cannon2024grid]. ∎
@@
We see that each such forest can, by adding k1k-1 edges of the components, be made into a spanning tree, and since every spanning tree can be split into $k$ pieces by arbitrarily removing k1k-1 edges, every spanning tree can be made this way. Since there are at most $n^{2}$ edges in $G$, the number of spanning trees is at most $n^{2(k-1)}$ times the number of such forests. Similarly, every such forest can be made by removing k1k-1 edges from a spanning tree, the number of such forests is at most $n^{2(k-1)}$ times the number of spanning trees (this bound can be reduced to $n^{k-1}$ by the structure of the tree). ∎
@@
By Lemma 2.1, each splittable tree uniquely defines a $k$-balanced spanning forest, so the number of them is at most the number of spanning trees. Further, each balanced tree can be defined by a balanced forest and k1k-1 edges, so there are at most $n^{2(k-1)}$ times as many spanning trees as balanced forests. ∎